### v1

# Electron Inertia and Magnetic Reconnection

Abstract
The finite electron mass can cause magnetic reconnection even in the absence of any other non-ideal effect in a magnetic evolution. It will be shown that when electron inertia is the only non-ideal effect in the evolution of the magnetic field $\vec{B}$, there is a related field that evolves ideally. This field is $\vec{\mathcal{B}}\equiv\vec{B}+\vec{\nabla}\times\left((m_{e}/ne^{2})\vec{j}\right)$ with $m_{e}$ the electron mass, $n$ the electron number density, and $\vec{j}$ the current density. Although the magnetic field is modified from its ideal evolution form by the electron inertia, the effect on particle trajectories, even electron trajectories, is small unless the current lies in thin sheets, which make $\vec{j}$ extremely large. The field $\vec{\mathcal{B}}$ is closely related to Voigt normalized magnetic field, which is defined by a Laplacian smoothing of $\vec{B}$. The difference between $\vec{\mathcal{B}}$ and $\vec{B}$ involves the relativistically invariant four-space Laplacian acting on $\vec{B}$ with a $c/\omega_{pe}$ smoothing distance; $\omega_{pe}$ is the plasma frequency.
I Introduction
The finite electron mass implies that a change in the current density in a plasma must be accompanied by an electric field, which gives a non-ideal evolution Biskamp ; Shivamoggi ; Andres:2014 . The concept is a close analogue to the London penetration depth of superconductors. The form the reconnection takes can be viewed as a physical example of a Voigt renormalization Constantin:2023 ; Yi-Min:2025 , $\vec{B}\rightarrow\vec{B}+(c/\omega_{pe})^{2}\nabla^{2}\vec{B}$, which smoothes the magnetic field over the scale of plasma skin depth $c/\omega_{pe}$.
A remarkable result is that the divergence-free modified magnetic field, $\vec{\mathcal{B}}$, which is of the form of Voigt renormalized magnetic field, has a perfect ideal evolution when the only non-ideal effect is the electron inertia. The field lines of $\vec{\mathcal{B}}$ have a velocity through space $\vec{u}_{\bot}$ and cannot break. This somewhat surprising result is of interest in itself but also has interesting implications for the guiding center drift equations. The actual effect on particle trajectories is small unless the current lies in thin sheets which make $\vec{j}$ extremely large.
II Modified magnetic field
In the frame of reference in which the ions are stationary, the current density $\vec{j}$ and the electron force balance equation are
$\vec{j}=-en\vec{v}_{e};$ (1)
$m_{e}n\frac{d\vec{v}_{e}}{dt}=-en(\vec{E}+\vec{v}_{e}\times\vec{B})$
$\hskip 50.58878pt-\vec{\nabla}\cdot\tensor{P}_{e}-m_{e}n\nu_{ei}\vec{v}_{e},\hskip 14.45377pt\mbox{so }$ (2)
$\vec{E}=\frac{\vec{j}\times\vec{B}}{en}-\frac{\vec{\nabla}\cdot\tensor{P}_{e}}{en}+\frac{m_{e}\nu_{ei}}{e^{2}n}\vec{j}+\frac{d}{dt}\left(\frac{m_{e}}{ne^{2}}\vec{j}\right).\hskip 14.45377pt$ (3)
The resistive term $\eta\vec{j}$ with $\eta\equiv m_{e}\nu_{ei}/(ne^{2})$ and the electron pressure/viscosity term $\vec{\nabla}\cdot\tensor{P}_{e}/en$ cause reconnection and are set to zero in order to study the effect of electron inertia.
Expanding $d\vec{v}_{e}/dt=\partial\vec{v}_{e}/\partial t+\vec{v}_{e}\cdot\vec{\nabla}\vec{v}_{e}$, reconnection will be studied when the electric field is
$\vec{E}=\frac{\vec{j}\times\vec{B}}{en}+\frac{\partial}{\partial t}\left(\frac{m_{e}}{ne^{2}}\vec{j}\right)-\frac{\vec{j}\cdot\vec{\nabla}}{en_{e}}\left(\frac{m_{e}}{ne^{2}}\vec{j}\right),\hskip 14.45377pt$ (4)
$\mbox{where}\hskip 14.45377pt\frac{m_{e}}{ne^{2}}\vec{j}=\frac{c^{2}}{\omega_{pe}^{2}}\mu_{0}\vec{j};$ (5)
$c/\omega_{pe}\approx 0.5$~cm when the electron density is $10^{20}$~m-3.
The evolution of the magnetic field is given by Faraday’s Law, $\partial\vec{B}/\partial t=-\vec{\nabla}\times\vec{E}$ so
$\frac{\partial}{\partial t}\left(\vec{B}+\vec{\nabla}\times\left(\frac{m_{e}}{ne^{2}}\vec{j}\right)\right)=$-$\vec{\nabla}\times\vec{\mathcal{E}}$ (6)
$\vec{\mathcal{E}}\equiv\frac{\vec{j}\times\vec{B}}{en}-\frac{\vec{j}\cdot\vec{\nabla}}{en}\left(\frac{m_{e}}{ne^{2}}\vec{j}\right).$ (7)
Although electron inertia breaks the lines of $\vec{B}$, it does not break the lines of
$\vec{\mathcal{B}}\equiv\vec{B}+\vec{\nabla}\times\left(\frac{m_{e}}{ne^{2}}\vec{j}\right)\hskip 14.45377pt\mbox{with }$ (8)
$\frac{\partial\vec{\mathcal{B}}}{\partial t}=-\vec{\nabla}\times\vec{\mathcal{E}}\hskip 14.45377pt\mbox{and }$ (9)
$\vec{\nabla}\cdot\vec{\mathcal{B}}=0.$ (10)
III Proof that $\vec{\mathcal{B}}$ has no reconnection
The proof has two parts. The first part is obtained by placing $\vec{\mathcal{E}}$ in a simplified form. The second is part is determining the condition for reconnection.
III.1 Simplied form for $\vec{\mathcal{E}}$
Equation (7) for $\vec{\mathcal{E}}$ can be rewritten in a simplified form
$\vec{\mathcal{E}}$ $=$ $\frac{\vec{j}\times\vec{\mathcal{B}}}{en}-\frac{e}{2m_{e}}\vec{\nabla}\left(\frac{m_{e}\vec{j}}{ne^{2}}\right)^{2}.$ (11)
The derivation of the simplified form starts with the Hall term.
$\frac{\vec{j}\times\vec{B}}{en}=\frac{\vec{j}\times\vec{\mathcal{B}}}{en}-\frac{\vec{j}}{en}\times(\vec{\mathcal{B}}-\vec{B}).$ (12)
The term
$\frac{m_{e}\vec{j}\times(\vec{\mathcal{B}}-\vec{B})}{ne^{2}}$
$\hskip 14.45377pt=\left(\frac{m_{e}\vec{j}}{ne^{2}}\right)\times\left(\vec{\nabla}\times\frac{m_{e}\vec{j}}{ne^{2}}\right)$ (13)
$\hskip 14.45377pt=\frac{1}{2}\vec{\nabla}\left(\frac{m_{e}\vec{j}}{ne^{2}}\right)^{2}-\left(\frac{m_{e}\vec{j}}{ne^{2}}\right)\cdot\vec{\nabla}\left(\frac{m_{e}\vec{j}}{ne^{2}}\right).\hskip 14.45377pt$ (14)
Consequently the Hall term has the form
$\frac{\vec{j}\times\vec{B}}{en}$ $=$ $\frac{\vec{j}\times\vec{\mathcal{B}}}{en}-\frac{e}{2m_{e}}\vec{\nabla}\left(\frac{m_{e}\vec{j}}{ne^{2}}\right)^{2}$ (15)
$+\left(\frac{\vec{j}}{ne}\right)\cdot\vec{\nabla}\left(\frac{m_{e}\vec{j}}{ne^{2}}\right),$
which implies Equation (7) for $\vec{\mathcal{E}}$ can be rewritten as Equation (11).
III.2 Reconnection condition
The proof that Equation (11) for $\vec{\mathcal{E}}$ implies that the lines of $\vec{\mathcal{B}}$ cannot reconnect starts with a purely mathematical equation, the representation of the field $\vec{\mathcal{E}}$ in terms of $\vec{\mathcal{B}}$.
Anywhere that $\vec{\mathcal{B}}\neq 0$, mathematics implies any vector $\vec{\mathcal{E}}$ in three-space can be represented as
$\vec{\mathcal{E}}+\vec{u}_{\bot}\times\vec{\mathcal{B}}=-\vec{\nabla}\Phi+\mathcal{E}_{r}\vec{\nabla}\ell,$ (16)
where $\ell$ is the the distance along a field line of $\vec{\mathcal{B}}$ with $\vec{\mathcal{B}}\cdot\vec{\nabla}\ell=\mathcal{B}$, and $\bot$ means perpendicular to $\vec{\mathcal{B}}$.
The potential $\Phi$ can always be chosen so $\mathcal{E}_{r}$ is constant along $\vec{\mathcal{B}}$. When $\mathcal{E}_{r}$ can be chosen to be zero the evolution of $\vec{\mathcal{B}}$ is ideal, and the field lines are carried with the velocity $\vec{u}_{\bot}$, the field line velocity Newcomb as shown below.
The validity of Equation (16) is proven by showing that all three components of $\vec{\mathcal{E}}$ can be represented. The parallel component is
$\frac{\vec{\mathcal{B}}}{\mathcal{B}}\cdot\vec$$\mathcal{E}$} $=$ $-\frac{\vec{\mathcal{B}}}{\mathcal{B}}\cdot\vec{\nabla}\Phi+\mathcal{E}_{r}$ (17)
$\frac{\vec{\mathcal{B}}}{\mathcal{B}}\cdot\vec{\nabla}\Phi$ $=$ $\left(\frac{\partial\Phi}{\partial\ell}\right)_{\alpha\beta},\hskip 14.45377pt\mbox{so }$ (18)
$\mathcal{E}_{||}$ $=$ $-\left(\frac{\partial\Phi}{\partial\ell}\right)_{\alpha\beta}+\mathcal{E}_{r}$ (19)
using Clebsch coordinates $(\alpha,\beta,\ell)$. Since $\vec{\mathcal{B}}$ is divergence free, it can be written in the Clebsch form $\vec{\mathcal{B}}=\vec{\nabla}\alpha\times\vec{\nabla}\beta$.
Equation (11) for $\vec{\mathcal{E}}$ then implies
$\frac{\partial\Phi}{\partial\ell}$ $=$ $\frac{e}{2m_{e}}\frac{\partial}{\partial\ell}\left(\frac{m_{e}\vec{j}}{ne^{2}}\right)^{2},$ (20)
which gives a single-valued expression for $\Phi$, so $\mathcal{E}_{r}$ can be taken to be zero. The perpendicular components of Equation (16) are fit by $\vec{u}_{\bot}$.
What is essentially Newcomb’s proof Newcomb that $\mathcal{E}_{r}=0$ ensures that the lines of $\vec{\mathcal{B}}$ are carried by the flow, $\vec{u}_{\bot}$, follows from Faraday’s Law. Using the Clebsch representation, $\partial\vec{\mathcal{B}}/\partial t=\vec{\nabla}\times((\partial\alpha/\partial t)\vec{\nabla}\beta-(\partial\beta/\partial t)\vec{\nabla}\alpha)$. Since $\vec{u}\times\vec{\mathcal{B}}=(\vec{u}_{\bot}\cdot\vec{\nabla}\beta)\vec{\nabla}\alpha-(\vec{u}_{\bot}\cdot\vec{\nabla}\alpha)\vec{\nabla}\beta$, when $\mathcal{E}_{r}=0$, the evolution is consistent with $d\alpha/dt\equiv\partial\alpha/\partial t+\vec{u}_{\bot}\cdot\vec{\nabla}\alpha=0$ and $d\beta/dt\equiv\partial\beta/\partial t+\vec{u}_{\bot}\cdot\vec{\nabla}\beta=0$. A field line is defined by giving $\alpha$ and $\beta$, and these labels of a field line are carried by the flow $\vec{u}_{\bot}$.
IV Implications
IV.1 Reconnection due to electron inertia
The purely mathematical Equation (16) applies to the magnetic field $\vec{B}$ as well as to $\vec{\mathcal{B}}$ once obvious notational changes have been made. The term $\mathcal{E}_{r}\vec{\nabla}\ell$ can be replaced in a torus by $(V_{\ell}/2\pi)\vec{\nabla}\varphi$ with $\varphi$ the toroidal angle. The loop voltage $V_{\ell}$ gives the topological changes in field.
The electric field is
$\vec{E}=\frac{\partial}{\partial t}\left(\frac{c^{2}}{\omega_{pe}^{2}}\mu_{0}\vec{j}\right)+\vec{\mathcal{E}}$ (21)
To lowest order in the short distance $c/\omega_{pe}$, the parallel component of the electric field is given by
$E_{||}$ $=$ $\frac{\vec{B}}{B}\cdot\frac{\partial}{\partial t}\left(\frac{c^{2}}{\omega_{pe}^{2}}\mu_{0}\vec{j}\right).$ (22)
The loop voltage on an irrational magnetic surface or along any of the closed magnetic field lines of a rational surface is
$V_{||}$ $=$ $\frac{1}{N_{t}}\int_{0}^{2\pi N_{t}}E_{||}\frac{B}{\vec{B}\cdot\vec{\nabla}\varphi}d\varphi\hskip 14.45377pt\mbox{as }N_{t}\rightarrow\infty\hskip 14.45377pt$ (23)
$=$ $\frac{1}{N_{t}}\int_{0}^{2\pi N_{t}}\frac{\vec{B}}{\vec{B}\cdot\vec{\nabla}\varphi}\cdot\frac{\partial}{\partial t}\left(\frac{c^{2}}{\omega_{pe}^{2}}\mu_{0}\vec{j}\right)d\varphi,$ (24)
which is not generally zero in an evolving magnetic field.
IV.2 Particle trajectories
The importance of magnetic field lines comes in large part from their effect on the trajectories of charged particles. When particles have a gyroradius $\rho\equiv v_{\bot}/\omega_{c}$ that is small compared to the size of the system, $a$, charged particles gyrate about $\vec{B}$ in near-circles of radius $\rho$. The velocity of the particle perpendicular to the magnetic field is $\vec{v}_{\bot}$, and the cyclotron frequency is $\omega_{c}\equiv qB/m$ with $q$ the charge and $m$ the mass of the particle.
When $|\rho/a|<<1$ and the magnetic and electric fields are slowly varying compared to $\omega_{c}$, the quantity $\mu\equiv mv_{\bot}^{2}/B$ is an accurately conserved adiabatic invariant. Averaging over the near-circular motion, the trajectory of a charged particle is accurately approximated by $d\vec{x}/dt=\vec{v}_{g}$, where $\vec{v}_{g}$ is known as the guiding center velocity. In time-independent magnetic and electric fields
$\vec{v}_{g}=\frac{v_{||}}{B}\vec{\nabla}\times(\vec{A}+\rho_{||}\vec{B})\hskip 14.45377pt\mbox{with }$ (25)
$\rho_{||}(H,\mu,\vec{x})\equiv\frac{v_{||}}{\omega_{c}}\hskip 14.45377pt\mbox{and }$ (26)
$H=\frac{1}{2}v_{||}^{2}+\mu B+q\Phi.$ (27)
The full equation was given by Boozer Drift:Boozer in 1980; the perpendicular part by Morozov and L. S. Solov’ev Morozov:1966 in 1966.
The vector potentials of $\vec{B}=\vec{\nabla}\times\vec{A}$ and $\vec{\mathcal{B}}=\vec{\nabla}\times\vec{\mathcal{A}}$ are related by
$\vec{\mathcal{A}}=\vec{A}+\frac{m_{e}}{ne^{2}}\vec{j}.$ (28)
Equation (25) for the guiding center velocity is then
$\vec{v}_{g}=\frac{v_{||}}{B}\vec{\nabla}\times\left(\vec{\mathcal{A}}+\rho_{||}\vec{B}-\frac{m_{e}}{ne^{2}}\vec{j}\right).$ (29)
This equation can be used to obtain the guiding center trajectories relative to the modified field, which has no reconnection due to electron inertia rather than relative of $\vec{B}$, which does reconnect due to electron inertia.
Comparing Equation (11) and Equation (16), the velocity $\vec{u}_{\bot}$ of the lines of $\vec{\mathcal{B}}$ is
$\vec{u}_{\bot}=\frac{\vec{\mathcal{B}}}{\mathcal{B}^{2}}\times\left(\frac{e}{2m_{e}}\vec{\nabla}\left(\frac{m_{e}\vec{j}}{ne^{2}}\right)^{2}-\frac{\vec{j}\times\vec{\mathcal{B}}}{en}\right).$ (30)
The calculation was done in the frame or reference in which the ions are stationary. The field lines could be taken to be stationary if the ions moved with a velocity $\vec{u}_{\bot}$. Nonetheless, the ratio of the perpendicular to the total current is
$\left|\frac{j_{\bot}}{j}\right|\approx\left(\frac{c}{\omega_{pe}}\right)^{2}\left|\frac{\mu_{0}j}{B}\right|\left|\vec{\nabla}\ln\left(\left|\frac{c^{2}}{\omega_{pe}^{2}}j\right|\right)\right|;$ (31)
$\left|\frac{\mu_{0}j}{B}\right|\approx\frac{1}{a}\frac{B_{\theta}}{B_{\varphi}}\approx\frac{1}{qR};$ (32)
$\left|\vec{\nabla}\ln\left(\left|\frac{c^{2}}{\omega_{pe}^{2}}j\right|\right)\right|\approx\frac{1}{a}.$ (33)
In a typical tokamak equilibrium, $a$ is the minor radius, $R\approx 3a$ is the major radius, and $q\approx 3$ is the safety factor. When the current density becomes extremely high by being concentrated in thin current sheets, the flow must be essentially parallel to the magnetic field lines to avoid $\vec{j}\times\vec{B}$ forces that cannot be balanced. As will be seen in Section IV.3, when $\left|(c/\omega_{pe})^{2}\nabla^{2}\vec{B}\right|<<\left|\vec{B}\right|$, the difference between $\vec{\mathcal{B}}$ and $\vec{B}$ is small. To lowest order in $c/\omega_{pe}$ and $\rho_{||}$, the guiding center velocity can be approximated as
$\vec{v}_{g}=\frac{v_{||}}{B}\vec{\nabla}\times\left(\vec{\mathcal{A}}+\left(\rho_{||}-\frac{c^{2}}{\omega_{pe}^{2}}\frac{\mu_{0}j_{||}}{B}\right)\vec{\mathcal{B}}\right).$ (34)
The ion gyroradius in a tokamak power plant will be comparable to $c/\omega_{pe}\approx 0.5$~centimeters and the electron gyroradius is about sixty times smaller. Since $qR\approx 20$~m, $(c/\omega_{pe})(\mu_{0}j_{||}/B)\approx 2.5\times 10^{-3}$, and the $\mu_{0}j_{||}/B$ term would seem to have a negligible effect on even the electron trajectories unless very thin current layers are formed.
IV.3 Voigt-regularized magnetic field
The relationship between $\vec{\mathcal{B}}$ and the Voigt-regularized magnetic field follows simply when the number density $n$ is constant in space and time. The generalized Ampere’s law is
$\mu_{0}\vec{j}=\vec{\nabla}\times\vec{B}-\mu_{0}\epsilon_{0}\frac{\partial\vec{E}}{\partial t},$ (35)
where $\mu_{0}\epsilon_{0}=1/c^{2}.$ When the spatial and temporal variation in the electron density are negligible, Equation (8) is
$\vec{\mathcal{B}}-\vec{B}$ $=$ $\frac{c^{2}}{\omega_{pe}^{2}}\vec{\nabla}\times\mu_{0}\vec{j}$ (36)
$=$ $\frac{c^{2}}{\omega_{pe}^{2}}\left(-\nabla^{2}\vec{B}-\frac{\partial}{\partial t}\vec{\nabla}\times\vec{E}\right)$ (37)
$=$ $-\frac{c^{2}}{\omega_{pe}^{2}}\Big(\nabla^{2}\vec{B}-\frac{1}{c^{2}}\frac{\partial^{2}\vec{B}}{\partial t^{2}}\Big)$ (38)
using Faraday’s Law. The relativistically invariant Laplacian is $\nabla^{2}\vec{B}-(1/c^{2})\partial^{2}\vec{B}/\partial t^{2}.$
Acknowledgements
This work received no external support.
Author Declarations
The author has no conflicts to disclose.
Data availability statement
Data sharing is not applicable to this article as no new data were created or analyzed in this study.